Mariko is paid a base salary plus commission for selling kitchen appliances. One week, her sales totaled $3800, and she earned $594. In a busier week, her sales totaled $5750 and she earned $652.50.
a) what rate of commission is Mariko paid?
b) what is her weekly base salary?
c) How much would she earn in a week if her sales totaled $4325?

Answer by mananth(16949)

Salary = x
commission = y
x+3800*0.01y=594
x+5750*0.01y=652.50
x+38y=594
x+57.5y=652.50
subtract the two
19.5y = 58.50
y= 58.50/19.5
y=3% the commission
plug value of y in first equation
x+3800*0.03=594
x+114=594
x=594-114
x=$480 the salary per week
..
480+0.03*4325=$609.75